The following information is provided:

Sales price per unit $480
Fixed costs $60,000
Break-even in units 2,000

The variable cost is:
Answer #1

Breakeven=Fixed expenses/Contribution margin

Contribution margin=(60,000/2000)=$30 per unit

Contribution margin=Sales-Variable cost

Variable cost=(480-30)

=$450 per unit